OpenStudy (anonymous):

-11x-10y+17 5x-7y+50 solve please?

OpenStudy (vanz):

y=-11/10x+17/10

OpenStudy (anonymous):

Eh? sorry, i don't understand this much.

OpenStudy (anonymous):

Is this a systems of equations?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Just one second and I'll have it

OpenStudy (anonymous):

wait a sec, where is the = sign?

OpenStudy (anonymous):

the second one should be 5x-7y=50, sorry.

OpenStudy (anonymous):

what about the first one? ...=17 ?

OpenStudy (anonymous):

yes, sorry again

OpenStudy (anonymous):

its alright, the shift key can get annoying sometimes

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

y = -5, x = 3

OpenStudy (anonymous):

thanks a lot. :D

OpenStudy (anonymous):

here is how i did this: I wanted to get x or y by itsself so i turned 5x - 7y = 50 into 5x= 7y + 50

OpenStudy (anonymous):

then, i divided by 5

OpenStudy (anonymous):

and got x = 7/5y + 10

OpenStudy (anonymous):

then i just plugged that whole thing into the x variable in the first equation like this -11(7/5y + 10) - 10y = 17, then you distribute the -11 and add together the two y terms and put the constant on the other side and get y by itsself. After that you can just plug y into either equation to get x